# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v1

@@ Abstract
Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data. Evaluated on three datasets, ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple quantitative and qualitative metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow
@@ 1 Introduction
Flow matching (Lipman et al., 2023) is an emerging paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions. Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available. In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment and perturbation responses (Wagner & Klein, 2020). Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data. Flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
@@
The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of complex cellular processes is affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
@@
Optimal transport (OT) has become a foundational framework to align spatially resolved samples and infer putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023). As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths to train velocity fields, thus overcoming the lack of generative capabilities in optimal transport. The OT formulation adopted in MOTFM, however, does not account for the contextual richness present in spatial transcriptomics and can result in trajectories that are statistically optimal yet biologically implausible (see Figure 3(a) in Appendix G.1 for an illustration). While recent studies have extended widely-used OT objectives (Halmos et al., 2025; Ceccarelli et al., 2025) for spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors for modeling temporal tissue dynamics (Figure 1). By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories. In summary, our contributions are as follows:
  • We leverage local tissue organization and local ligand–receptor communication patterns to extract biologically meaningful features from spatial omics data, and encode them into a biologically-informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
  • We design two novel integration schemes—cost-based and entropy-based—that incorporate the prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
  • Comprehensive experiments on regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline methods under both interpolation and extrapolation settings across metrics that capture biological plausibility and statistical fidelity (Section 4).
@@
Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data. Prior knowledge acts as a soft filter that discourages implausible transitions while preserving flexibility in trajectory inference.
@@ 2.1 Flow Matching Basics
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani et al., 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
@@ 2.2 Conditional Flow Matching
The central idea of conditional flow matching is to express the target probability path $p_{t}$ via a mixture of more manageable conditional probability paths (Lipman et al., 2023). By marginalizing over some conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},\quad u_{t}(\bm{x})=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (2)
where $q(\bm{z})$ denotes the distribution of the conditioning variable $\bm{z}$, and $p_{t}(\bm{x}|\bm{z})$ is selected such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$. To avoid the intractable integrals, Lipman et al. (2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original flow matching objective in terms of gradient computation:
@@
By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thus avoiding the intractable integration operation. Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields:
$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{z}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (4)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ stands for the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ stand for the corresponding derivative. In particular, Lipman et al. (2023) set $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};0,I)$ to a Gaussian distribution with mean $x_{1}$ and standard deviation $\sigma$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be verified. Tong et al. (2024) further generalized the application scope to arbitrary source distributions, by setting
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (5)
This choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. Based on Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
@@ 2.3 Flow Matching with Optimal Transport Couplings
The conditionals construction specified by Equation 5 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani et al., 2008). Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
Since the classical Kantorovich’s formulation (refer Appendix E) has computational complexity that is cubic with respect to the sample size, a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
$\pi^{*}_{\mathrm{eot}}(\epsilon):={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (6)
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. The optimization problem in Equation 6 can be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013). Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 16 in Appendix E for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.

### v2

@@ Abstract
Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data. Comprehensive experiments show that ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow.
Machine Learning, ICML
@@ 1 Introduction
Flow matching (Lipman et al., 2023) is an emerging generative paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions. Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many application domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available. In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment, and perturbation responses (Wagner and Klein, 2020). Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data. In contrast, flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
Table 1: Comparisons of state-of-the-art spatiotemporal OT and flow matching methods. In particular, ContextFlow excels at efficiently inferring biologically plausible trajectories by incorporating spatial priors into minibatch-OT-based flow-matching frameworks.
Method Generative Prior Knowledge Unified Velocity Field OT Runtime
DeST-OT (Halmos et al., 2025) $O(N^{3})$
TOAST (Ceccarelli et al., 2025) $O(N^{3})$
PASTE (Zeira et al., 2022) $O(N^{3})$
CFM (Lipman et al., 2022)
MOTFM (Tong et al., 2024) $O(N^{2})$
ContextFlow (ours) $O(N^{2})$
@@
The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of complex cellular processes are affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
@@
Optimal transport (OT) has become a foundational framework for aligning spatially resolved samples and inferring putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023). As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths for training velocity fields, thereby overcoming the lack of generative capabilities in optimal transport. The OT formulation adopted in MOTFM, however, does not account for the contextual richness of spatial transcriptomics and can yield trajectories that are statistically optimal yet biologically implausible (see Figure 4(a) in Appendix G.1 for an illustration). While recent studies (Halmos et al., 2025; Ceccarelli et al., 2025) have extended OT objectives to spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
@@
Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data. Prior knowledge acts as a soft filter, discouraging implausible transitions while preserving flexibility in trajectory inference.
To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors to model temporal tissue dynamics (Figure 1). By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories. Table 1 highlights the key advantages of ContextFlow, compared with the aforementioned existing methods, where detailed discussions on related works are postponed to Appendix A.
In summary, our contributions are as follows:
  • We leverage the structure of local tissue organization and ligand–receptor communication to extract biologically meaningful features from spatial omics data and encode them into an informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
  • We design two novel integration schemes—cost-based and entropy-based—that incorporate prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
  • Comprehensive experiments on three regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline flow matching methods under both interpolation and extrapolation settings across multiple evaluation metrics that capture biological plausibility and statistical fidelity (Section 4).
@@ 2.1 Flow Matching Basics
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani and others, 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
@@ 2.2 Conditional Flow Matching
The central idea of conditional flow matching is to express the target probability path via a mixture of more manageable conditional probability paths (Lipman et al., 2023). By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},$ (2)
$u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$
where $q(\bm{z})$ stands for the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
To avoid the intractable integrals, (Lipman et al., 2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original FM objective in terms of gradient computation:
@@
By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thereby avoiding the intractable integration operation. Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields, defined as follows:
$p_{t}(\bm{x}|\bm{z})$ $=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),$ (4)
$u_{t}(\bm{x}|\bm{z})$ $=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (5)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative.
In particular, Lipman et al. (2023) selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated. Tong et al. (2024) further generalized the application scope to arbitrary source distributions by setting:
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\sigma_{t}(\bm{z})=\sigma.$ (6)
Note that the choice in Equation 6 satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. According to Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
@@ 2.3 Flow Matching with Optimal Transport Couplings
The conditionals construction specified by Equation 6 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani and others, 2008). Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
Since the classical Kantorovich’s formulation has computational complexity that is cubic in terms of the sample size (see Appendix B.1 for its formal definition), a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
${\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}$ $\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})$ (7)
$\qquad\qquad+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. For simplicity, we use $\pi^{*}_{\mathrm{eot}}(\epsilon)$ to denote the optimal solution to the minimization problem in Equation 7. Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013). Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 14 in Appendix B.1 for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.